Question 2. (20 pts.) a) Simplify 2 < (57/2)) (2 < (--/3) 2« (-%) b) Solve Z + Z2 + 1 = 0
Answer #1

simplify: 57 2 As we know that we can write = cosot isino , we get Using this (२०) (२०१४) Hedis.५ - 2x4 .83 (54-7 8 (154 - 476) solve z +2²+1=0 ₂4 +2²+1=0 We can write above equation in such a forum (2²+ y)² - z² = 0 [(22+4)- z] [zº+1+2) (22-2+1) (24 22+42+4=0 422+42+1 +3=0 (22+1 )2-(UBI) = 0 (2249-vzi) (Qz+) = 0 VEL 2= -1 + Pence, the complete solution for value of zis -I hope it will help you.

Thanks!